If p is a parameter of positive numbers, show that all members of the family of ellipses x2/ (a2 + p) + y2 / (b2 + p) = 1 have the same foci.


1
Expert's answer
2021-08-18T12:39:20-0400

If the equation of an ellipse is "\frac{x^2}{m^2}+\frac{y^2}{n^2}=1," the the foci are "(-c,0)" and "(c,0)," where "c=\sqrt{m^2-n^2}." In our case,


"c=\sqrt{m^2-n^2}=\sqrt{(\sqrt{a^2+p})^2-(\sqrt{b^2+p})^2}\\
=\sqrt{a^2+p-(b^2+p)}=\sqrt{a^2-b^2}."


We conclude that the foci are "(-\sqrt{a^2-b^2},0)" and "(\sqrt{a^2-b^2},0)." Since the foci do not depend on the parameter "p," the family of ellipses "\frac{x^2}{a^2 + p} + \frac{y^2}{b^2 + p} = 1" have the same foci.